Grade 12Physical Chemistry

calculate the bond order of [He--H]+ from molecular orbital theory ….......

Ankur Saha
9 Years agoGrade 12

2 Answers

Vikas TU
9 Years ago
First calculate the total number of electrons in it that is => 2 + 1 – 1 => 2
Therfore the elcrtocic config. from MT concept would be => 1sigmas2
2 electrons are in BMO and 0 electrons are in ABMO.
Hence the bond order becomes => (2-0)/2 = 1 
Jack
5 Years ago
Total  electron  2+1-1=2
electronic configuration of molecular orbital ∂1s2 so bonding electron =2.antibonding electron =0 so bond order = (2-0)÷2 =1